Find an equation of the line through the given points. Write the equation in standard form.
Through (-1,1) and (1, -13).

Answer by tommyt3rd(5050)

m=(-13-1)/(1-(-1))=-14/2=-7
y-1=-7(x-(-1))
y-1=-7x-7
7x+y=-6